OpenStudy (anonymous):

Linear algebra help please! Determine the values of the parameter (alpha) the following planes in R^3 are perpendicular: (alpha)x+10y-8z=2 (alpha)x+(alpha)y-2z=-11

OpenStudy (anonymous):

If their normal vectors are orthogonal, the planes are as well.

OpenStudy (anonymous):

\[ \begin{bmatrix} \alpha \\ 10 \\-8 \end{bmatrix} \cdot \begin{bmatrix} \alpha \\ \alpha \\-2 \end{bmatrix} = 0 \]

OpenStudy (anonymous):

Great!!! Thank you so much! I'd give you an eggroll if you were here :D